OpenStudy (anonymous):

Find the remainder of the division problem shown below. (x⁴+x³-3x-3)÷(x+2)

OpenStudy (anonymous):

if you only want the remainder, replace x by -2 in the numerator. that is in \[x^4+x^3-3x-3\] and see what you get. that will be the remainder

OpenStudy (anonymous):

I got -21/x+2 ???

OpenStudy (anonymous):

Note the formula: P(x)/D(x)= Q(x)+R(x)/D(x), where R(x) is our remainder, P(x) is the polynomial and Q(x) is our quotient and D(x), the divisor. Here i got R(x)/D(x)=11/(x+2), this implies that our R(x), the remainder is 11.

OpenStudy (precal):